Alpha’s balance-of-payments data for 2016 are shown below. All figures are in billions of dollars.

Category Value
Goods exports $+ 50
Goods imports – 35
Service exports + 25
Service imports – 10
Net investment income – 5
Net transfers + 10
Balance on capital account 0
Foreign purchases of Alpha assets + 25
Alpha purchases of assets abroad – 60

Instructions: Input all amounts as positive values.

a. What is the balance on goods?

  

b. What is the balance on goods and services?

  

c. What is the balance on the current account?

    

d. What is the balance on the capital and financial account?

  

Suppose Alpha sold $15 billion of official reserves abroad to balance the capital and financial account with the current account.

e. Does Alpha have a balance-of-payments deficit or does it have a surplus?
Answer #1

Ans 1.) (a) Balance on good :

= Export of goods - Import of goods

= 50 - 35

Balance on goods = $15 billion

(b) Balance on goods and service :

= > Export of goods and service - Import of goods and service

=> (50 + 25 ) - (35 + 10)

= > 75 - 45

Balance on goods and service = $30 billion

(c) Balance on current account:

=> Net trade balance + Net transfers + Net investment income

=> 30 - 5 + 10

Current account balance = $35 billion

(d) Balance on capital account/financial account :

=> Capital account + financial account

Financial account:

Sale of domestic assets - purchase of foreign assets:

=> 25 - 60 = -$35 billion

Balance on capital account is calculated as :

= 0 + (-35) = -$35 billion